Find the lump sum deposit which will give $9,430 in 10 years in an account paying 1.24% compounded monthly. 


1
Expert's answer
2021-07-14T10:11:26-0400

First, convert R as a percent to r as a decimal

"r = \frac{1.24}{100}\\

r = 0.0124" rate per year,

"A = P(1 +\frac{ r}{n})^{nt}"

A = Accrued amount,P is the Principal ,r is annual nominal interest rate as a decimal,n is number of compounding periods per unit of time and t time in decimal years

"9430 = P\times(1 + \frac{0.0124}{12})^{(12)(10)}\\

9430 = P\times(1.001033333)^{(120)}\\
9430 =1.131943399P\\P=\frac{9430} {1.131943399}\\P=\$8330.81"